Two cells each of end 1.5 volt and internal resistance 5 ohm are joined in series with a resistance box and a resistance coil of 10 ohm. What resistance must be unplugged in the box in order that that p.d. across the ends of the ohm coil may be 0.1 volt? 

Grade:12th pass

1 Answers

Shivansh
5 years ago
3V total emf, 10 ohm internal resistance, 0.1volt is dropped across the box, i e, 2.9V is across 10ohm.
current = 2.9/10 = 0.29A

hencer resistance = 0.1/0.29 = 0.345 ohm